Question 12 (1 point) Charles just borrowed $3,565.00 and agreed to repay it with three equal payments. The first will be in 2 months, the second in 4 months and the third in 6 months Usin g a simple interest rate of 5.25% per annum, calculate the size of each payment. today as the focal date and express your answer to 2 decimal places Use Your Answer Answer
Answer #1

We have following information -

Borrowed Amount = $ 3,565

Simple interest rate = 5.25% pa

thus, Simple interest rate for a month = 0.0525/12 = 0.004375

Repayment = 3 equal payment in 2nd, 4th & 6th month.

Let's assume repayment value is $ X

The Present value(PV) of all three equal payment at simple interest should be equal to Borrowed amount today.

Thus,

Borrowed Amount = PV of repayments

Repayment 1 rt Borrowuedamount 7l3565 = 1(0.0043752) 1(0.0043754(0.004375 6)

1.00875 1.0175 1.02625

1.00875 1.0175 1.02625

3565- X (0.9913 0.9828 0.9744)

3565 = X (2.9485)

3565 -2.9489

X 1209.09 -

Thus, the equal repayment is $ 1209.09

Please note - PV of equal repayment is calculated based on Simple interest factor.